# DreamAudio: Customized Text-to-Audio Generation with Diffusion Models

### v1

@@ Abstract
With the development of large-scale diffusion-based and language-modeling-based generative models, impressive progress has been achieved in text-to-audio generation. Despite producing high-quality outputs, existing text-to-audio models mainly aim to generate semantically aligned sound and fall short on precisely controlling fine-grained acoustic characteristics of specific sounds. As a result, users that need specific sound content may find it challenging to generate the desired audio clips. In this paper, we present DreamAudio for customized text-to-audio generation (CTTA). Specifically, we introduce a new framework that is designed to enable the model to identify auditory information from user-provided reference concepts for audio generation. Given a few reference audio samples containing personalized audio events, our system can generate new audio samples that include these specific events. In addition, two types of datasets are developed for training and testing the customized systems. The experiments show that the proposed model, DreamAudio, generates audio samples that are highly consistent with the customized audio features and aligned well with the input text prompts. Furthermore, DreamAudio offers comparable performance in general text-to-audio tasks. We also provide a human-involved dataset containing audio events from real-world CTTA cases as the benchmark for customized generation tasks. Demos are available at https://yyua8222.github.io/DreamAudio_demopage/.
Index Terms:
audio generation, diffusion model, retrieval argumentation, customized generation, AIGC
@@ I Introduction
Audio generation, as a crucial technology for enabling artificial intelligence generated content (AIGC) [1], has gained significant interest from the research community. In recent years, conditional audio generation has become a popular paradigm, where music, speech, or general sound effects are generated based on a variety of conditions, such as text [2], images [3, 4], and videos [5, 6]. This opened new opportunities for a range of potential applications, including audio generation for movies [7], games [8], and audiobooks [9].
@@
Facilitated by advances in diffusion-based generative models [10, 11] and large-scale audio-language datasets [12, 13, 14, 15], several models have been developed for audio generation with text prompts as conditions, such as AudioLDM [2], AudioGen [16], DiffSound [17], TANGO [18], Make-an-Audio2 [19], Re-AudioLDM [20], and AudioLDM2 [21]. Building on the “semantic prior” learned from large collections of datasets [22] to associate textual concepts with audio features, these models have shown strong capabilities in generating audio samples of high quality, fidelity and diversity. For example, when provided with the text prompt dog barking, the model leverages this particular “semantic prior” to generate various audio clips depicting dog barking across varying species, emotions, and durations.
@@
Despite significant progress in diffusion-based generative systems [23, 24, 25, 26, 27], current text-to-audio generation systems often lack the flexibility to customize content based on personalized intentions, particularly when dealing with rare or unseen sounds. This can cause problems in real-world applications for multimedia content production which often requires the generation of audio samples to be tailored for specific features. For example, it can be challenging for current TTA models to generate sounds that are consistent with the text prompt “a monster is fighting with a Minion”. This is because such audio events, like “monster fight” and “Minion talk” are unique and rare, or with a specific timbre, which can hardly be found from any current training data. Hence, users often need to go through several rounds of trial and error to produce the desired output and may encounter difficulties in achieving the optimal result [28].
@@
Figure 1: A demo of DreamAudio for audio generation with customized content of “monster fighting” and “Minion talking”. The system takes both the text prompt and user-provided audio-caption pairs as the reference concepts, and generate audio content consistent with the description ”Monster is fighting with a Minion”.
@@
To address this challenge, several works have explored the ideas for generating rare or unseen audio events, especially for few-shot or zero-shot scenarios. One of the pioneering works is the Re-AudioLDM [20], which is a retrieval-based technique to improve the performance for the generation of rare audio events using audio-caption pairs retrieved from external data. Similarly, AudioBox TTA-RAG [29] introduced a retrieval embedding module to improve the performance for few-shot and zero-shot audio generation tasks. Although these previous approaches effectively improve the semantic accuracy in generating infrequent or unseen audio events, they are still unable to provide explicit control over the sound effects generated, as specified by users.
@@
In this paper, we propose a new TTA task, namely, customized text-to-audio generation (CTTA), where the content produced by the generation system can be controlled and customized by users. For instance, the system is capable to generate the dog barking sound from a specific dog with unique timber, or the sound for a specific monster which is not included in the traning dataset. We refer to these featured audio events as user reference concepts. To achieve this, we introduce DreamAudio, a latent diffusion-based system with flow matching for the CTTA task. More specifically, inspired by ControlNet [30], we propose a multi-reference customization (MRC) structure for the generator module to control the content generated by the model. We design a new group of encoder blocks to extract features from the user-provided reference concepts. Then, the system is trained to fuse such customized features for the generation of the audio output. With a new cross-attention module linking the provided audio concepts with their corresponding language descriptions, our proposed method provides a training-free pipeline that is capable of generating audio samples, consistent with the reference concepts. Figure 1 shows an example of CTTA, where the model is enabled to generate the audio sample with customized “monster fighting” and “Minion talking”, in terms of the provided reference concepts.
@@
To facilitate model development and evaluation, we create two datasets with different formats by concatenating and overlapping different audio events, called Customized-Concatenation and Customized-Overlay. In addition, we collect various special audio events for customized cases, such as fiction and comic artificial sounds, and create a small-scale dataset that better aligns with real-world multimedia production scenarios as a benchmark for CTTA. We perform experiments on these two datasets, and the results show that our method achieves state-of-the-art (SOTA) results in terms of audio generation with customized content. Furthermore, DreamAudio presents comparable results to previous SOTA models on general text-to-audio generation tasks using AudioCaps [31], illustrating substantial advancements in customized TTA tasks and competitive performance in general TTA tasks.
@@
  • We propose a multi-reference customization structure, which allows the features from the reference audio to be fused for text-to-audio generation.
  • We develop two new datasets for training and evaluating the proposed method and establish a new benchmark for content-customized text-to-audio generation from real-world audio events.
  • Our experiments show that DreamAudio significantly enhances the ability of TTA models for customized audio generation and achieves state-of-the-art CTTA performance.
@@ II-A Diffusion Models
Diffusion-based models [11, 32] have demonstrated improved performance in generative tasks for image [33, 24, 34, 35], audio [36, 37], and video [38, 39]. In the realm of audio synthesis, researchers initially followed the design used in image generation, and adapted it for audio generation, based on mel-spectrogram [40, 41] and waveform [42, 43, 44]. Due to the involvement of high-dimensional data [45, 46], the generation systems based on waveform and spectrogram representations can be difficult to train and slow to infer. To address this limitation, recent diffusion-based models work in a latent space through encoding pipelines [26]. Specifically, current SOTA systems follow an encoder-decoder framework [47], trained to generate the target features in a latent space, which are then decoded and reconstructed into waveforms through vocoders [48, 49, 50].
@@ II-B Conditional Audio Generation
Controllable generation has emerged as an important area in audio synthesis, enabling users to guide generative models with specific constraints or conditions to produce desired outcomes [30]. Text-driven audio generation has gained significant attention. AudioGen [16] employs a conditional language modeling pipeline that generates audio waveforms directly. AudioLDM [2] employs the Contrastive Language-Audio Pretraining (CLAP) [51] model to derive the embeddings of audio and text, and the use them as conditions to guide the training and inference of the latent diffusion model or audio generation. Specifically, Audio is trained to generate feature representations of target audio within the latent space, followed by a variational autoencoder (VAE) decoder to reconstruct the spectrogram from the latent representation, finally using a vocoder to reconstruct the waveforms from the spectrogram. Make-an-Audio develops a pseudo-prompt enhancement strategy to generate extra audio-caption pairs with large-scale compositions to alleviate the data scarcity problem. Tango [52] uses a structure similar to Audio, but replaces the CLAP model with a FLAN-T5 [53].
@@
Due to the diversity of the training dataset [12, 31, 13], these previous methods allow the model to generate highly diverse samples for the same prompt. However, these models are not designed for CTTA tasks and often struggle to generate content from user preferences [20], which is the problem to be addressed in this paper.
@@
Figure 2: The inference pipeline of the DreamAudio. The input prompt and reference concept are encoded in two paralleled paths through the Flan-T5 Encoder and the reference audio feature is encoded by the VAE Encoder. Along with the noised data $\boldsymbol{z}_{\lambda}$, four inputs are forwarded to the generator with MRC structure to generate the denoised data, followed by the VAE decoder and vocoder to reconstruct the final output.
II-C Customized Content Generation
@@
Current methods for customized content generation are developed dominantly for images, such as [54, 55, 56, 57, 58, 59, 60]. They could be categorized into training-based customization [22, 61, 62, 63], and tailoring-based customization [64, 65, 66, 67, 68, 56, 69]. In training-based methods, the generation models are fine-tuned with the embeddings of specific subjects, in order to bind a unique identifier and control the generated contents. In tailoring-based methods, customization is achieved by integrating the target feature with the generated feature during the inference stages to minimize the fine-tuning processes.
@@
The CTTA task remains to be explored. Most similar to the CTTA task are retrieval-based TTA tasks, developed recently to improve the performance of TTA models in the few-shot and zero-shot cases, such as Re-AudioLDM [20], which incorporates external audio features into the generation process using retrieval-augmented approaches [70, 71]. This system demonstrates more stable performance on low-occurrence audio events in TTA generation. Building on this concept, AudioBox TTA-RAG [29] introduces a retrieval-information embedding module to enhance the capability in zero-shot generation. Although these systems can improve the TTA generation performance in the few-shot and zero-shot cases, the quality of the audio samples is limited by the database for audio retrieval. Furthermore, these models lack the capability to generate specific content for customized tasks and the research for CTTA tasks remains to be explored.
@@
In this paper, we design a novel CTTA system, which allows the generated audio content to be tailored and controlled in terms of user preference derived from the reference concepts, in addition to the text prompts. The ability to achieve customization not only enhances the creative process, but also improves controllability in TTA systems, thereby addressing a crucial practical challenge for deploying TTA systems in real-world applications.
@@ III-A Overview of the Proposed Model
Given only a few audio-language reference pairs $\{(\mathbf{a}_{1},\mathbf{t}_{1}),(\mathbf{a}_{2},\mathbf{t}_{2}),\dots,(\mathbf{a}_{K},\mathbf{t}_{K})\}$, where each reference audio clip $\mathbf{a}_{k}$ is associated with a reference caption $\mathbf{t}_{k}$, k=1,,Kk=1,...,K, the CTTA task aims to generate the target audio $\mathbf{a}$ conditioned on the input textual prompt $\boldsymbol{c}$ and the reference concepts, derived from pairs of reference audios $\{\mathbf{a}_{1},...,\mathbf{a}_{K}\}$ with their corresponding captions $\{\mathbf{t}_{1},...,\mathbf{t}_{K}\}$.
To address this task, we propose the DreamAudio system, as illustrated in Figure 2. The system is composed of several modules, including the text encoder (e.g. Flan-T5) to obtain the embeddings of the input text prompt and reference captions, and the audio encoder (e.g. VAE encoder) to obtain the embeddings of the reference audios, which are followed by the feature generator built on rectified flow matching (RFM). To customize the audio features with reference concepts, multi-reference customization (MRC) structure is designed for the feature generator. The customized audio embeddings are then converted to spectrograms using a VAE decoder, which is then turned into a waveform using a vocoder (e.g. BigVGAN [50]).
We first discuss the calculation of text and audio embeddings in Section III-B. Then, we present the implementation of our framework for diffusion-based feature generatorin Section III-C, starting with the preliminaries of module training strategy, followed by the details of the proposed MRC architecture. Finally, Section III-D introduces the process for reconstructing the target audio feature.

### v2

@@ Abstract
With the development of large-scale diffusion-based and language-modeling-based generative models, impressive progress has been achieved in text-to-audio generation. Despite producing high-quality outputs, existing text-to-audio models mainly aim to generate semantically aligned sound and fall short of controlling fine-grained acoustic characteristics of specific sounds. As a result, users who need specific sound content may find it difficult to generate the desired audio clips. In this paper, we present DreamAudio for customized text-to-audio generation (CTTA). Specifically, we introduce a new framework that is designed to enable the model to identify auditory information from user-provided reference concepts for audio generation. Given a few reference audio samples containing personalized audio events, our system can generate new audio samples that include these specific events. In addition, two types of datasets are developed for training and testing the proposed systems. The experiments show that DreamAudio generates audio samples that are highly consistent with the customized audio features and aligned well with the input text prompts. Furthermore, DreamAudio offers comparable performance in general text-to-audio tasks. We also provide a human-involved dataset containing audio events from real-world CTTA cases as the benchmark for customized generation tasks. Demos are available at https://yyua8222.github.io/DreamAudio_demopage/.
@@ I Introduction
Audio generation, as a crucial technology for enabling artificial intelligence generated content (AIGC) [4], has gained significant interest from the research community. In recent years, conditional audio generation has become a popular paradigm, where music, speech, or general sound effects are generated based on a variety of conditions, such as text [43], images [77, 70], and videos [54, 29]. This opened up new opportunities for a range of potential applications, including audio generation for movies [37], games [18], and audiobooks [49].
@@
Facilitated by advances in diffusion-based generative models [74, 26, 80] and large-scale audio-language datasets [21, 53, 76, 88], several models have been developed for audio generation with text prompts as conditions, such as AudioLDM [43], AudioGen [38], DiffSound [85], TANGO [22], Make-an-Audio2 [27], Re-AudioLDM [91], and AudioLDM2 [44]. Building on the “semantic prior” learned from large collections of datasets [64] to associate textual concepts with audio features, these models have shown strong capabilities in generating audio samples of high quality, fidelity and diversity. For example, when provided with the text prompt dog barking, the model leverages this particular “semantic prior” to generate various audio clips depicting dog barking across varying species, emotions, and durations.
@@
Despite significant progress in diffusion-based methods [62, 61, 2, 63, 14], current text-to-audio generation systems often lack the flexibility to customize content based on personalized intentions. This can cause problems in real-world multimedia production, which often requires the generation of audio samples to be tailored for specific features. For example, it can be challenging for current TTA models to generate sounds that are consistent with the text prompt “a monster is fighting with a Minion”. This is because audio events such as “monster fight” and “Minion talk” are unique and rare, or with a specific timbre, which can hardly be found in any current training data. Hence, users often need to go through multiple rounds of trial and error to produce the desired output and may encounter difficulties in achieving the optimal result [68].
@@
Figure 1: An illustration of DreamAudio for audio generation with customized content of “monster fighting” and “Minion talking”. The system takes both the text prompt and user-provided audio-caption pairs as the reference concepts, and generate audio content consistent with the description “Monster is fighting with a Minion”.
@@
To address this challenge, several works have explored the ideas for generating rare or unseen audio events, especially for few-shot or zero-shot scenarios. One of the pioneering works is the Re-AudioLDM [91], which applies retrieval-based techniques to improve the performance for the generation of rare audio events using audio-caption pairs retrieved from external data. Similarly, AudioBox TTA-RAG [86] introduced a retrieval embedding module to improve the performance for few-shot and zero-shot audio generation tasks. Although these previous approaches effectively improve the semantic accuracy in generating infrequent or unseen audio events, they are still unable to provide explicit control over the sound effects generated, as specified by users.
@@
In this paper, we propose a new TTA task, namely, customized text-to-audio generation (CTTA), where the audio content produced by the generation system can be controlled and customized by users. For instance, the system is capable to generate the dog barking sound from a specific dog with unique timber, or the sound for a specific monster fighting which is not included in the traning dataset. We refer to these featured audio events as user reference concepts. To achieve this, we introduce DreamAudio, a latent diffusion-based system with flow matching for the CTTA task. Inspired by ControlNet [93], we propose a multi-reference customization (MRC) structure for the generator module to identify the reference concepts and control the generated content. More specifically, we design a new group of encoder blocks to extract features from the user-provided reference concepts. The system is trained to fuse such customized features for the generation of the audio output. By incorporating multiple cross-attention modules that link the target text prompt with the corresponding reference text, our proposed approach establishes a pipeline capable of generating user-preferred audio samples without requiring concept-specific fine-tuning during inference. In contrast to tuning-based customization methods, which necessitate model updates for each new reference even after the model has been fully trained, DreamAudio directly extracts and integrates features from the new references within a single forward pass. Figure 1 shows an example of CTTA, where the model is enabled to generate the audio sample with customized “monster fighting” and “Minion talking”, in terms of the provided reference concepts.
@@
To facilitate model development and evaluation, we create two datasets with different formats by concatenating and overlapping different audio events, called Customized-Concatenation and Customized-Overlay. In addition, we collect several special audio events and then manually design various customized cases, creating a small-scale dataset that more closely reflects real-world multimedia production scenarios and serves as a benchmark for CTTA. Experiments conducted on these datasets demonstrate that our method effectively handles customized audio generation and establishes a strong baseline for this new CTTA task.
@@
  • We propose a multi-reference customization (MRC) structure, which allows the features from the reference audio to be fused with input prompts for text-to-audio generation.
  • We develop two new datasets for model training and evaluation, and establish a new benchmark for content-customized text-to-audio generation.
  • Our experiments show that DreamAudio significantly enhances the ability of TTA models for customized audio generation and establishes a strong baseline for this new CTTA task.
@@ II-A Diffusion Models
Diffusion-based models [26, 75] have demonstrated improved performance in generative tasks for image [12, 61, 66, 67], audio [5, 36, 46], and video [73, 25]. In the realm of audio synthesis, researchers initially followed the design used in image generation, and adapted it for audio generation based on mel-spectrogram [58, 10] and waveform [40, 41, 9]. Due to the involvement of high-dimensional data [35, 51], the representations for waveform and spectrogram can be difficult to train and slow to infer. To address this limitation, recent diffusion-based models work in a latent space through encoding pipelines [63]. Specifically, current state-of-the-art (SOTA) systems follow an encoder-decoder framework [90], trained to generate the target features in a latent space, which are then decoded into waveforms through vocoders [33, 39, 42].
@@ II-B Conditional Audio Generation
Controllable generation has emerged as an important area in audio synthesis, enabling users to guide generative models with specific constraints or conditions to produce desired outcomes [93]. Text-driven audio generation has gained significant attention. AudioGen [38] employs a conditional language modeling pipeline that generates audio waveforms directly. AudioLDM [43] employs the Contrastive Language-Audio Pretraining (CLAP) [84] model to generate the embeddings of audio and text. These embeddings are then used as conditions to guide the training and inference of the latent diffusion model for audio generation. Specifically, AudioLDM is trained to generate feature representations of target audio within the latent space, followed by a variational autoencoder (VAE) decoder to reconstruct the spectrogram from the latent representation. Make-an-Audio [27] develops a pseudo-prompt enhancement strategy to generate extra audio-caption pairs with large-scale compositions to alleviate the data scarcity problem. Tango [22] uses a structure similar to AudioLDM, but replaces the CLAP model with Flan-T5 [60].
@@
Due to the diversity of the training dataset [21, 32, 53], these previous methods allow the model to generate highly diverse samples for the same prompt. However, these models are not designed for CTTA tasks and often struggle to generate content in terms of user preferences [91], which is the problem to be addressed in this paper.
@@
Figure 2: The inference pipeline of the DreamAudio. The input prompt and reference concept are encoded in two paralleled paths through the Flan-T5 Encoder and the reference audio feature is encoded by the VAE Encoder. Along with the noisy data $\boldsymbol{z}_{\lambda}$, four inputs are forwarded to the generator with the MRC structure to generate the denoised data, followed by the VAE decoder and vocoder to reconstruct the final output waveform.
II-C Audio Generation with Flow Matching
Flow matching has recently emerged as an efficient alternative to diffusion-based generative models. By learning a deterministic flow path between noise and target distribution [92], flow matching significantly reduces the number of inference steps required for high-quality synthesis, leading to improved efficiency and enhanced performance. FlashAudio [47] is the first audio generation model to employ flow matching, providing fast and high-fidelity results. LAFMA [23] further exploits flow-matching models and reduces the number of inference steps to ten without compromising performance. More recently, TangoFlux [28] and Stable Audio [16] have applied flow matching to large-scale text-to-audio models, achieving SOTA performance on several audio-generation benchmarks. These studies demonstrate that flow matching provides an efficient and scalable framework for audio synthesis, enabling faster sampling and strong performance. Our work builds upon this framework and applies flow matching as the core backbone.
II-D Customized Generation
@@
Current methods for customized content generation are developed dominantly for images[24, 89, 3, 7, 20, 31, 78]. Customized image generation could be categorized into tuning-based customization [64, 19, 81, 1], and tuning-free customization [50, 65, 83, 3, 13, 11]. In tuning-based methods, such as DreamBooth [64], the generation models are fine-tuned with the embeddings of specific subjects to control the generated contents. In tuning-free methods, such as Freecustom [13], customization is achieved by integrating the target feature with the generated feature during the inference stages to minimize the fine-tuning processes. Our DreamAudio resembles tuning-free customization methods. While the model itself is trained from scratch to learn general customization capabilities, it does not require any additional adaptation or parameter updates when encountering unseen user-provided reference concepts during inference.
II-D1 Speech and Music Generation
Several customized systems have been explored in speech and music generation. For speech generation, ViT-TTS [45] introduces a visual-text encoder that extracts additional visual scene information from reference images to improve text to speech (TTS) generation performance. F5-TTS [8] incorporates reference speech as conditioning signals to control speaking style, prosody, and voice characteristics. In the music domain, Plitsis et al. [57] adapted user-specific concepts from DreamBooth [64] and fine-tune the text embeddings for personalized music style control. Unlike CTTA systems, these existing methods mainly adjust global acoustic attributes such as speaker style, prosody, timbre, genre, or overall musical texture. Their conditioning signals depend on complete speech or full musical phrases, whereas CTTA requires fine-grained, event-level controllability, making it fundamentally different from these style-based generation tasks.
II-D2 General Audio Generation
FreeAudio [30] and TG-Diff [15] are two works designed to customize the timing of sound events in general audio generation. These two models both applied tuning-free customization strategies. More specifically, TG-Diff achieves temporal controllability by learning per-second embeddings, while FreeAudio achieves customization by applying a strategy for decoupling and aggregating attention tokens to fill-in the time-window tokens. However, both systems focus on timing control in text-to-audio generation, which differs from our goal on controlling the acoustic audio events. In addition, these two methods focus on training the time-related embeddings and tokens to match the timing window. Such requirement does not involve zero-shot scenarios, as most temporal patterns can be covered by massive training data with different temporal features.
II-E Retrieval-Based Generation
@@
Another task similar to CTTA is retrieval-based TTA, which was developed recently to improve the performance of TTA models in the few-shot and zero-shot cases. An example is Re-AudioLDM [91], which incorporates external audio features into the generation process using retrieval-augmented approaches [71, 6]. This system demonstrates more stable performance on low-occurrence audio events in TTA generation. Building on this concept, AudioBox TTA-RAG [86] introduces a retrieval-information embedding module to enhance the capability in zero-shot generation. Although these systems can improve the TTA generation performance in the few-shot and zero-shot cases, the quality of the audio samples is limited by the database for audio retrieval. Furthermore, these models lack the capability to generate specific content for customization tasks, and research for CTTA tasks remains to be explored.
@@
In this paper, we design a novel CTTA system, which allows the generated audio content to be tailored and controlled in terms of user preference derived from the reference concepts. The ability to achieve customization not only enhances the creative process, but also improves controllability in the event level in TTA systems. Hence, we aim to address a crucial practical challenge for deploying TTA systems in real-world applications.
@@ III-A Overview of the Proposed Model
Given only a set of k=1,,Kk=1,...,K audio-language reference pairs $\{(\mathbf{a}_{1},\mathbf{t}_{1}),(\mathbf{a}_{2},\mathbf{t}_{2}),\dots,(\mathbf{a}_{K},\mathbf{t}_{K})\}$, where each reference audio clip $\mathbf{a}_{k}$ is associated with a reference caption $\mathbf{t}_{k}$, the CTTA task aims to generate the target audio $\mathbf{a}$ conditioned on the input textual prompt $\boldsymbol{c}$ and the reference concepts, derived from pairs of reference audios $\{\mathbf{a}_{1},...,\mathbf{a}_{K}\}$ with their corresponding captions $\{\mathbf{t}_{1},...,\mathbf{t}_{K}\}$.
To address this task, we propose the DreamAudio system, as illustrated in Figure 2. The system is composed of several modules, including the text encoder (e.g. Flan-T5 [60]) to obtain the embeddings of the input text prompt and reference captions, and the audio encoder (e.g. the variational autoencoder (VAE) as built in [43]) to obtain the embeddings of the reference audios, which are followed by the feature generator built on rectified flow matching (RFM) [48]. To customize the audio features with reference concepts, a multi-reference customization (MRC) structure is designed for the feature generator. The customized audio embeddings are then converted to spectrograms using a VAE decoder, which are then turned into waveforms using a vocoder (e.g. BigVGAN [42]).
In the remainder of this section, we first discuss the calculation of text and audio embeddings. Then, we present our framework for diffusion-based feature generator, starting with the preliminaries of the module training strategy, followed by the details of the proposed MRC architecture. Finally, we introduce the process for reconstructing the target audio.